OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)2 – 16? Identify the x–intercepts by graphing the quadratic equation. x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –16) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, –16) x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –15) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, –15)

OpenStudy (anonymous):

To find the x-intercept's, set y equal to 0. 0 = (x - 1)^2 - 16 16 = (x - 1)^2 (x - 1) = ±4 x = -3 or x = 5 So (-3, 0) and (5, 0) are coordinates of the x-intercepts. To find the y-intercepts, set x equal to 0. y = (0 - 1)^2 - 16 y = 1 - 16 y = -15

OpenStudy (anonymous):

Wait is that 2 a regular 2 or an exponent ^2?

OpenStudy (anonymous):

I think the answer is D

OpenStudy (anonymous):

an exponent 2

OpenStudy (anonymous):

ok then its D for sure

OpenStudy (anonymous):

thank you brandy!!!

OpenStudy (anonymous):

Welcome lol